Find the roots of (-1+sqrt 3i)^3/2 leaving the answer in polar from.
Thanks

Answer by stanbon(75887)

Find the roots of (-1+sqrt 3i)^3/2 leaving the answer in polar from.
-----
Polar form:: r(cos(t)+i*sin(t))
----------------------
r = sqrt[1^2 + (sqrt(3))^2] = sqrt[4] = 2
-----------------------
t = arctan(sqrt(3)/(-1)] = arctan(-sqrt(3)) = (2/3)pi
-------------------
Ans: r^(3/2)(cos[(2pi/3+2npi)/(3/2) + i*sin[(2pi/3+2npi)/(3/2)]
----------------
If n = 0, Ans = r^(3/2)[cos[(4/9)pi + i*sin[(4/9)pi]
-----
If n = 1, Ans = r^(3/2)[cos[(4/9)pi+(4/3)pi] + isin[(4/9)pi + (4/3)pi]
---
= r^(3/2)*cis[(16/9)pi]
--------------
Cheers,
Stan H.